OpenStudy (idealist10):

Solve y'=(-x+3y-14)/(x+y-2) implicitly.

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (anonymous):

Have you tried applying that method we discussed last time?

OpenStudy (idealist10):

u=x+h v=y+k dy/dx=dv/du dv/du=(-(u-h)+3(v-k)-14)/((u-h)+(v-k)-2) dv/du=(-u+3v+(h-3k-14))/(u+v+(-h-k-2)) h-3k-14=0 -h-k-2=0 k=-4 h=2 dv/du=(-u+3v)/(u+v) dv/du=(-1+3v/u)/(1+v/u) v=uz dv/du=u*dz/du+z d*dz/du+z=(-1+3z)/(1+z) u*dz/du=(-1+3z)/(1+z)-z u*dz/du=(-1+3z)/(1+z)-(z(1+z))/(1+z) =(-1+3z-z-z^2)/(1+z) =(-1+2z-z^2)/(1+z) =(z^2-2z+1)/(z+1) =(z-1)^2/(z+1) (z+1)/(z-1)^2 dz=du/u u=z-1 du=dz z+1=u+2 ln abs(z-1)-2/(z-1) ln abs(z-1)-2/(z-1)=ln abs(u)+C ln abs(z-1)-ln abs(u)-2/(z-1)=C ln abs((z-1)/u)-2/(z-1)=C am I right so far?

OpenStudy (idealist10):

Thank you, sithandgiggles.